OpenStudy (anonymous):

Last month, Mayfield Paper Company sold 58 more than 3 times as many reams of paper as Superior Paper Products. If the two companies sold a total of sold a total of 870 reams of paper, how many reams of paper did the Mayfield Paper Company sell? Please also include work. thanks

OpenStudy (dumbcow):

a+b = 870 a = 3b+58 By substitution (3b+58) +b = 870 4b = 870 -58 b = 203 a = 870 - 203 = 667

OpenStudy (dumbcow):

a is Mayfield paper co. b is superior paper

OpenStudy (anonymous):

thanks